OpenStudy (anonymous):

Seriously HELP ME.. For f(x) = 1/x+5 and g(x)= x^2+2 Find (f o g)(x) (g o f)(6)

OpenStudy (anonymous):

I just really don't understand

OpenStudy (anonymous):

for (f o g)(x) you take the g(x) and you substitute that equation in for x in the f(x) equation if that makes sense?

OpenStudy (anonymous):

Soo... no not really

OpenStudy (anonymous):

lol alright so you take the second one, which is the g in this case. and the g(x) is x^2+2. You then take the x^2+2 and plug it in for x in the f(x) formula

OpenStudy (anonymous):

oohhh that's all?

OpenStudy (anonymous):

yup

OpenStudy (khaakenajaf):

when finf fog put g(x) in f(x) f( x^2+2) = 1/ (x^2+2)+5 g(1/x+5)= [1/x+5]^2+2

OpenStudy (khaakenajaf):

got it

OpenStudy (anonymous):

oohh wow It was that simple.. I'm getting all frustrated over here

OpenStudy (khaakenajaf):

no prob

OpenStudy (anonymous):

can you help me out some more? i'm gonna put another question up.